# Gravitational and electromagnetic Cherenkov radiation constraints in modified dispersion relations

### v1

@@ 2.3 Threshold condition
Let us start discussing the case $\mathbb{A},\mathbb{B}>0$, where the threshold momentum is the minimum value of $k$ for the process to be allowed. Then, $k_{\text{min}}>p$ if $\mathbb{B}<\mathbb{A}p$ and the process is forbidden, as can be seen from equation (18). Otherwise, the process is allowed with minimum momentum given by equation (21); in particular, the process is allowed for all $k$ when $\mathbb{B}>2\mathbb{A}p$, see equation (17). The analysis is reversed when considering $\mathbb{A},\mathbb{B}<0$. Now equation (21) establishes the maximum momentum up to where the process is allowed. As long as $\mathbb{B}<2\mathbb{A}p$ the process is permitted, and for $\mathbb{B}<\mathbb{A}p$ it is possible for all values of $k$; for $\mathbb{B}>2\mathbb{A}p$ Cherenkov emission is not possible as $k_{\text{max}}<0$. A similar analysis can be carried out solving equation (14) for different values of $\alpha$ and $\beta$, and study whether the process is allowed or not attending to the signs of $\mathbb{A}$ and $\mathbb{B}$ and to equations (17) and (18). Table 1 illustrates the solutions to the threshold condition in the massless limit for the values $\alpha,\beta\in\{0,1,2\}$.
@@ 3 Dynamics of generalised Cherekov radiation
In this section we focus on the calculation of the rate of energy loss $\text{d}E/\text{d}t$ of the particle $a$ in the GCR process $a\to a+b$. We present the general procedure for the tree-level diagram shown in figure 1 for an arbitrary interaction vertex. Then we study the particular cases where the emitted particle is either a photon $\gamma$ (electromagnetic CR) or a graviton $h$ (gravitational CR).
@@ 3.1 Decay rate and energy loss
It is well known peskin1995introduction that the differential decay rate of a two body process in an arbitrary reference frame in a LI scenario is given by
@@
It is important to understand whether equation (22) is allowed to be used in a LI breaking scenario, as is the case with our MDRs. When LI is not broken, the canonical commutation relations between creation and annihilation operators of the field are $[a_{\bm{p}},a_{\bm{p}^{\prime}}^{\dagger}]=(2\pi)^{3}\delta^{(3)}(\bm{p}-\bm{p}^{\prime})$ when the usual factor $(2E_{\bm{p}})^{-1/2}$ in the momentum integral of the field operators is included. This factor on the field operators is reflected in the structure of the LIPS in equation (22). This is no longer true when LI does not hold. Indeed, the wave function of the particles involved in the process will have a different normalization condition, including corrections due to the LI breaking term in the MDR. Hence, to ensure the canonical normalization condition, field operators will no longer be normalized by a factor $(2E_{\bm{p}})^{-1/2}$ and the LIPS in equation (22) should receive additional corrections. Nevertheless, as we shall see during this section, the matrix elements $\mathcal{M}$ that we will consider violate LI, and so no corrections must be taken into account coming from the LIPS as long as we compute the decay rate in the preferred frame. Equation (22) is then valid for the purposes of this work Jacobson_2006 .
Let us now compute the integrals in equation (22). First of all, if we use the conservation of the $3-$momentum we can immediately integrate over $\bm{p}^{\prime}$ and obtain
@@
where $\mathcal{M}$ is assumed to be evaluated at $p^{\prime}=(E_{\bm{p}-\bm{k}},\bm{p}-\bm{k})$. To integrate over $\bm{k}$ we can use spherical coordinates and fix the $3-$momentum $\bm{p}$ along the $z$ axis. Note that this is possible since the MDR is invariant under rotations (and thus is the decay rate). As will be seen, the matrix element $\mathcal{M}(p\to p^{\prime},k)$ for the processes we consider does not depend on the azimutal angle, so
@@
When deriving the threshold condition in section 2, it can be shown from equation (11) that up to order $\mathcal{O}(\Theta_{c})$
@@
which substituting in equation (25) gives
@@
Note that we have approximated $E_{\bm{p}}=p+\mathcal{O}(\mathbb{A})$ in the denominator for the same reason why we did not consider higher order terms arising from the LIPS. The integration limits for $k$ are determined by the threshold condition (14), being $k=0$ and $k=p$ the lowest and highest integration limits possible, respectively. $H(\Theta_{c})$ is the Heaviside function and establishes whether the process is kinematically allowed or forbidden.
Equation (28) is valid for any process $a\to a+b$ with the MDR proposed with a matrix element $\mathcal{M}$ that does not depend on the azimutal angle. However, the emission rate of the $b$ particles does not allow us to impose any constraint on the LIVPs, unlike the energy loss rate of the $a$ particle. To consider the energy carried off by the particle $b$, we must insert the energy $E_{\bm{k}}=k+\mathcal{O}(\mathbb{B})$ of this particle in the integral of equation (28), obtaining the following energy loss rate of the particle $a$:
@@
To evaluate the matrix element $\mathcal{M}$ we will consider, for simplicity, the case where the particle $a$ is a complex scalar field, even though the constraints obtained in section 4 for the LIVPs are applied to fermionic particles. It is expected that spin corrections are of order $\mathcal{O}(1)$ and, therefore, are not significant enough to modify the order of magnitude of the constraints Moore_2001 .
@@ 3.2 Electromagnetic Cherenkov radiation
Let us first consider the process described in figure 1 where the particle $b$ is a photon $\gamma$. Feynman rules for a charged complex scalar field applied to this process give the following matrix element schwartz2013quantum :
@@
Here $\epsilon_{\mu}$ is the polarization vector of the photon, and $e_{a}$ the electric charge of the particle $a$. The analysis we will carry out does not distinguish between polarizations, hence we must sum over the two transverse polarization states of the photon when computing the squared matrix element. Using the conservation of $3-$momentum we get
@@
The physical polarizations are perpendicular to the $4-$momentum of the photon, therefore $k^{\mu}\epsilon_{\mu}(k)=0$. The product $p^{\mu}\epsilon_{\mu}(k)$ is easily calculated taking into account that the two polarizations are perpendicular to each other and the angle between $\bm{p}$ and $\bm{k}$ is $\theta_{c}$. This yields
@@
where we have used equation (12), considered the LI violating terms up to first order, and again denoted $p=|\bm{p}|$. Substituting this matrix element in equation (29), the electromagnetic energy loss rate is then given by
@@
where the integration limits $k_{\text{min}}$ and $k_{\text{max}}$ are obtained by solving the threshold condition (14). Also, note that if $\Theta_{c}(k)<0$ for all $k\in[0,p]$ the Heaviside function in equation (33) imposes the energy loss rate to be equal to zero, as Cherenkov emission is kinematically forbidden.
@@ 3.3 Gravitational Cherenkov radiation
We now consider the case where the particle $b$ in figure 1 is a graviton $h$. The Feynman rules for the process $a\to a+h$, with $a$ a complex scalar field can be found in reference Han_1999 :
@@
Here $G_{\text{N}}$ is the gravitational constant, and $\epsilon^{\mu\nu}$ is the polarization tensor for the graviton field $h_{\mu\nu}$. This tensor is traceless ($\epsilon_{\mu\nu}\eta^{\mu\nu}=0$) and transverse ($\epsilon_{\mu\nu}k^{\mu}=0$), and can be constructed in terms of the polarization vectors of massive vector bosons Han_1999 . As in the electromagnetic case, we do not distinguish between the polarizations of the graviton and thus we sum over the two physical polarization states. Using equation (36) and substituting in the matrix element (35) one finds
@@
Comparing this expression with that corresponding to the electromagnetic CR process, equation (32), it can be noted that for the gravitational case instead of the electric charge $e_{a}^{2}$ we have the gravitational constant times the $3-$momentum squared $G_{\text{N}}p^{2}$, and the emission is reduced for small opening angles, having $\sin^{4}{\theta_{c}}$ rather than $\sin^{2}{\theta_{c}}$. Substituting the matrix element (37) in the energy loss rate (29) and considering the lowest order in $\Theta_{c}$ in equation (12) we obtain
@@
We first notice that the energy loss rate (38) of the vacuum gravitational CR, due to the tensor nature of the gravitational field $h$, is of order $\mathcal{O}(\Theta_{c}^{2})$, whereas the energy loss rate (33) of the electromagnetic CR is of order $\mathcal{O}(\Theta_{c})$ because of the vector nature of the EM field. On the other hand, the electromagnetic coupling constant is much stronger than the gravitational coupling for energies much lower than the Planck scale: the factor $e_{a}^{2}/(2\pi)$ in equation (33) is typically of order $\mathcal{O}(10^{-2})$, while in contrast the factor $G_{\text{N}}p^{2}$ in (38) is of order $\mathcal{O}(10^{-14})$ for the highest energy particles observed in the universe. Therefore, as expected, the energy loss rate due to gravitational CR is much smaller than that caused by the emission of electromagnetic CR. So, for gravitational CR the decay rate of a particle of astrophysical origin can be comparable to its travel time and, therefore, it must be taken into account in order to constrain the LIVPs. Unlike for the electromagnetic CR, a kinematic analysis is not enough to determine an upper bound for $\mathbb{A}$ and $\mathbb{B}$.
@@
The indefinite integral in equation (38) can be computed in the general case $m\neq 0$ and $\alpha,\beta\geq 0$ in terms of hypergeometric functions, although it is not particularly illuminating. In addition, we also need to calculate the threshold momentum $k_{\text{th}}$ for the integration limits, which cannot be done analytically for arbitrary values of $\alpha$ and $\beta$. Nevertheless, the energy loss rate of the $a$ particle can be obtained for particular values of the parameters. For example, taking $\mathbb{A}=0$, one gets in the massless limit
@@
as it was obtained in reference Kiyota_2015  [^3] This result corresponds to equation (13) of Kiyota_2015 taking $\delta=0$ and identifying our exponent $\beta$ with $\alpha-2$ (do not confuse this exponent with that of the particle $a$) and our parameter $\mathbb{B}$ with $-A/2$. for $k_{\text{min}}=0$ and $k_{\text{max}}=p$. This result reduces to that presented in reference Moore_2001 if we additionally impose $\beta=0$.
@@ 4 Constraints on the Lorentz invariance violating parameters
We are now interested in obtaining estimations for the LIVPs $\mathbb{A}$ and $\mathbb{B}$ appearing in $n_{a}$ and $n_{b}$ using the energy loss rate of vacuum gravitational CR (38). We shall use high energy astrophysical observations for this purpose, arguing that UHECRs offer the most stringent bounds. We will consider different values of $\alpha$ and $\beta$ to discuss the corresponding phenomena and illustrate it in figure 2 for $\alpha=2$, $\beta=0$ and $\alpha=1$, $\beta=2$.
@@ 4.1 Constraints from the energy loss rate
In section 3 we discussed that the decay rate of a particle of astrophysical origin due to gravitational CR is expected to be of the same order as the time of propagation, which implies that the energy loss rate has to be taken into account in order to estimate the LIVPs of both the particle $a$ and the theory of gravity. Computing the integral in equation (38) and solving the differential equation would give the maximum time travel $t$ possible for a given momentum $p$ in terms of $\mathbb{A}$, $\mathbb{B}$ Moore_2001 ; Kimura_2012 . Thus, constraints on $\mathbb{A}$ and $\mathbb{B}$ can be imposed using high energy astrophysical observations if the distance travelled $ct$ and the momentum $p$ are known.
@@
To illustrate how constraints can be estimated, let us consider the simple case where $\alpha=\beta=0$. Attending to equations (17) and (18) we see that the process is allowed when $\mathbb{B}\geq\mathbb{A}$. Note that this condition is equivalent to the particle $a$ having a greater phase velocity than the particle $b$, $c_{a}>c_{b}$. Integration of equation (38) in the massless limit is straightforward and yields
@@
Here we see that constraints are more stringent for high energy particles that have travelled a long distance $ct$; note, however, that the energy of the particle plays a more relevant role since the dependence is of the form $(tp^{3})^{-1/2}$. Another interesting feature is that the constraint is only imposed for the difference between the LIVPs, so one of them remains unfixed: as long as this difference is small enough, the values of $\mathbb{A}$ and $\mathbb{B}$ can be arbitrarily large. One should be careful with this last statement as our results have been derived assuming that $|\mathbb{A}p^{\alpha}|$ and $|\mathbb{B}k^{\beta}|$ are very small compared to unity. We have checked that both of these features, the dependence on $p$ and $t$ and the impossibility of fixing both $\mathbb{A}$ and $\mathbb{B}$ simultaneously, are present for different values of $\alpha$ and $\beta$.
It is also worth to mention the case where $\mathbb{A}=0$ and $\beta$ is a positive real number. Performing a similar calculation in the massless limit $m=0$ shows that damping from gravitational CR is not relevant if
@@
which corresponds to the result derived in reference Kiyota_2015 . The constraint on $\mathbb{B}^{(\beta)}p^{\beta}$ exhibits the same dependence on the energy and the time travel as equation (41), and we see that the value of $\mathbb{B}^{(\beta)}$ decreases for higher values of $\beta$. This has been used to derive much more stringent constraints in the gravitational sector using the absence of vacuum CR for $\beta\gtrsim 0$ rather than using direct detection of GWs, in particular the events GW150914 and GW151226 Yunes_2016 . The same behaviour is observed when the dispersion relation of $a$ particles is modified, that is, when $\mathbb{A}\neq 0$.

### v2

@@ 2.3 Threshold condition
Let us start discussing the case $\mathbb{A},\mathbb{B}>0$, where the threshold momentum is the minimum value of $k$ for the process to be allowed. Then, $k_{\text{min}}>p$ if $\mathbb{B}<\mathbb{A}p$ and the process is forbidden, as can be seen from Eq. (18). Otherwise, the process is allowed with minimum momentum given by Eq. (21); in particular, the process is allowed for all $k$ when $\mathbb{B}>2\mathbb{A}p$; see Eq. (17). The analysis is reversed when considering $\mathbb{A},\mathbb{B}<0$. Now Eq. (21) establishes the maximum momentum up to where the process is allowed. As long as $\mathbb{B}<2\mathbb{A}p$ the process is permitted, and for $\mathbb{B}<\mathbb{A}p$ it is possible for all values of $k$; for $\mathbb{B}>2\mathbb{A}p$ Cherenkov emission is not possible as $k_{\text{max}}<0$. A similar analysis can be carried out solving Eq. (14) for different values of $\alpha$ and $\beta$, and study whether the process is allowed or not attending to the signs of $\mathbb{A}$ and $\mathbb{B}$ and to Eqs. (17) and (18). Table 1 illustrates the solutions to the threshold condition in the massless limit for the values $\alpha,\beta\in\{0,1,2\}$.
@@ 3 Dynamics of generalised Cherekov radiation
In this Sec. we focus on the calculation of the rate of energy loss $\text{d}E/\text{d}t$ of the particle $a$ in the GCR process $a\to a+b$. We present the general procedure for the tree-level diagram shown in Fig. 1 for an arbitrary interaction vertex. Then we study the particular cases where the emitted particle is either a photon $\gamma$ (electromagnetic CR) or a graviton $h$ (gravitational CR).
@@ 3.1 Decay rate and energy loss
It is well known peskin1995introduction that the differential decay rate of a two body process in an arbitrary Ref. frame in a LI scenario is given by
@@
It is important to understand whether Eq. (22) is allowed to be used in a LI breaking scenario, as is the case with our MDRs. When LI is not broken, the canonical commutation relations between creation and annihilation operators of the field are $[a_{\bm{p}},a_{\bm{p}^{\prime}}^{\dagger}]=(2\pi)^{3}\delta^{(3)}(\bm{p}-\bm{p}^{\prime})$ when the usual factor $(2E_{\bm{p}})^{-1/2}$ in the momentum integral of the field operators is included. This factor on the field operators is reflected in the structure of the LIPS in Eq. (22). This is no longer true when LI does not hold. Indeed, the wave function of the particles involved in the process will have a different normalization condition, including corrections due to the LI breaking term in the MDR. Hence, to ensure the canonical normalization condition, field operators will no longer be normalized by a factor $(2E_{\bm{p}})^{-1/2}$ and the LIPS in Eq. (22) should receive additional corrections. Nevertheless, as we shall see during this Sec., the matrix elements $\mathcal{M}$ that we will consider violate LI, and so no corrections must be taken into account coming from the LIPS as long as we compute the decay rate in the preferred frame. Eq. (22) is then valid for the purposes of this work Jacobson_2006 .
Let us now compute the integrals in Eq. (22). First of all, if we use the conservation of the 3-momentum we can immediately integrate over $\bm{p}^{\prime}$ and obtain
@@
where $\mathcal{M}$ is assumed to be evaluated at $p^{\prime}=(E_{\bm{p}-\bm{k}},\bm{p}-\bm{k})$. To integrate over $\bm{k}$ we can use spherical coordinates and fix the 3-momentum $\bm{p}$ along the $z$ axis. Note that this is possible since the MDR is invariant under rotations (and thus is the decay rate). As will be seen, the matrix element $\mathcal{M}(p\to p^{\prime},k)$ for the processes we consider does not depend on the azimuthal angle, so
@@
When deriving the threshold condition in Sec. 2, it can be shown from Eq. (11) that up to order $\mathcal{O}(\Theta_{c})$
@@
which, substituting in Eq. (25), gives
@@
Note that we have approximated $E_{\bm{p}}=p+\mathcal{O}(\mathbb{A})$ in the denominator for the same reason why we did not consider higher order terms arising from the LIPS. The integration limits for $k$ are determined by the threshold condition (14), where $k=0$ and $k=p$ the lowest and highest integration limits possible, respectively. $H(\Theta_{c})$ is the Heaviside function and establishes whether the process is kinematically allowed or forbidden.
Eq. (28) is valid for any process $a\to a+b$ with the MDR proposed with a matrix element $\mathcal{M}$ that does not depend on the azimuthal angle. However, the emission rate of the $b$ particles does not allow us to impose any constraint on the LIVPs, unlike the energy loss rate of the $a$ particle. To consider the energy carried off by the particle $b$, we must insert the energy $E_{\bm{k}}=k+\mathcal{O}(\mathbb{B})$ of this particle in the integral of Eq. (28), obtaining the following energy loss rate of the particle $a$:
@@
To evaluate the matrix element $\mathcal{M}$ we will consider, for simplicity, the case where the particle $a$ is a complex scalar field, even though the constraints obtained in Sec. 4 for the LIVPs are applied to fermionic particles. It is expected that spin corrections are of order $\mathcal{O}(1)$ and, therefore, are not significant enough to modify the order of magnitude of the constraints Moore_2001 .
@@ 3.2 Electromagnetic Cherenkov radiation
Let us first consider the process described in Fig. 1 where the particle $b$ is a photon $\gamma$. Feynman rules for a charged complex scalar field applied to this process give the following matrix element schwartz2013quantum :
@@
Here $\epsilon_{\mu}$ is the polarization vector of the photon, and $e_{a}$ the electric charge of the particle $a$. The analysis we will carry out does not distinguish between polarizations; hence, we must sum over the two transverse polarization states of the photon when computing the squared matrix element. Using the conservation of 3-momentum we get
@@
The physical polarizations are perpendicular to the 4-momentum of the photon and therefore $k^{\mu}\epsilon_{\mu}(k)=0$. The product $p^{\mu}\epsilon_{\mu}(k)$ is easily calculated taking into account that the two polarizations are perpendicular to each other and the angle between $\bm{p}$ and $\bm{k}$ is $\theta_{c}$. This yields
@@
where we have used Eq. (12), considered the LI violating terms up to first order, and again denoted $p=|\bm{p}|$. Substituting this matrix element in Eq. (29), the electromagnetic energy loss rate is then given by
@@
where the integration limits $k_{\text{min}}$ and $k_{\text{max}}$ are obtained by solving the threshold condition (14). Also, note that if $\Theta_{c}(k)<0$ for all $k\in[0,p]$ the Heaviside function in Eq. (33) imposes the energy loss rate to be equal to zero, as Cherenkov emission is kinematically forbidden.
@@ 3.3 Gravitational Cherenkov radiation
We now consider the case where the particle $b$ in Fig. 1 is a graviton $h$. The Feynman rules for the process $a\to a+h$, with $a$ a complex scalar field can be found in Ref. Han_1999 :
@@
Here $G_{\text{N}}$ is the gravitational constant, and $\epsilon^{\mu\nu}$ is the polarization tensor for the graviton field $h_{\mu\nu}$. This tensor is traceless ($\epsilon_{\mu\nu}\eta^{\mu\nu}=0$) and transverse ($\epsilon_{\mu\nu}k^{\mu}=0$), and can be constructed in terms of the polarization vectors of massive vector bosons Han_1999 . As in the electromagnetic case, we do not distinguish between the polarizations of the graviton and thus we sum over the two physical polarization states. Using Eq. (36) and substituting in the matrix element (35) one finds
@@
Comparing this expression with that corresponding to the electromagnetic CR process, Eq. (32), it can be noted that for the gravitational case instead of the electric charge $e_{a}^{2}$ we have the gravitational constant times the 3-momentum squared $G_{\text{N}}p^{2}$, and the emission is reduced for small opening angles, having $\sin^{4}{\theta_{c}}$ rather than $\sin^{2}{\theta_{c}}$. Substituting the matrix element (37) in the energy loss rate (29) and considering the lowest order in $\Theta_{c}$ in Eq. (12) we obtain
@@
We first notice that the energy loss rate (38) of the vacuum gravitational CR, due to the tensor nature of the gravitational field $h$, is of order $\mathcal{O}(\Theta_{c}^{2})$, whereas the energy loss rate (33) of the electromagnetic CR is of order $\mathcal{O}(\Theta_{c})$ because of the vector nature of the EM field. On the other hand, the electromagnetic coupling constant is much stronger than the gravitational coupling for energies much lower than the Planck scale: the factor $e_{a}^{2}/(2\pi)$ in Eq. (33) is typically of order $\mathcal{O}(10^{-2})$, while in contrast the factor $G_{\text{N}}p^{2}$ in (38) is of order $\mathcal{O}(10^{-14})$ for the highest energy particles observed in the Universe. Therefore, as expected, the energy loss rate due to gravitational CR is much smaller than that caused by the emission of electromagnetic CR. So, for gravitational CR the decay rate of a particle of astrophysical origin can be comparable to its travel time and, therefore, it must be taken into account in order to constrain the LIVPs. Unlike for the electromagnetic CR, a kinematic analysis is not enough to determine an upper bound for $\mathbb{A}$ and $\mathbb{B}$.
@@
The indefinite integral in Eq. (38) can be computed in the general case $m\neq 0$ and $\alpha,\beta\geq 0$ in terms of hypergeometric functions, although it is not particularly illuminating. In addition, we also need to calculate the threshold momentum $k_{\text{th}}$ for the integration limits, which cannot be done analytically for arbitrary values of $\alpha$ and $\beta$. Nevertheless, the energy loss rate of the $a$ particle can be obtained for particular values of the parameters. For example, taking $\mathbb{A}=0$, one gets in the massless limit
@@
as it was obtained in Ref. Kiyota_2015  [^3] This result corresponds to Eq. (13) of Kiyota_2015 taking $\delta=0$ and identifying our exponent $\beta$ with $\alpha-2$ (do not confuse this exponent with that of the particle $a$) and our parameter $\mathbb{B}$ with $-A/2$. for $k_{\text{min}}=0$ and $k_{\text{max}}=p$. This result reduces to that presented in Ref. Moore_2001 if we additionally impose $\beta=0$.
@@ 4 Constraints on the Lorentz invariance violating parameters
We are now interested in obtaining estimations for the LIVPs $\mathbb{A}$ and $\mathbb{B}$ appearing in $n_{a}$ and $n_{b}$ using the energy loss rate of vacuum gravitational CR (38). We shall use high energy astrophysical observations for this purpose, arguing that UHECRs offer the most stringent bounds. We will consider different values of $\alpha$ and $\beta$ to discuss the corresponding phenomena and illustrate it in Fig. 2 for $\alpha=2$, $\beta=0$ and $\alpha=1$, $\beta=2$.
@@ 4.1 Constraints from the energy loss rate
In Sec. 3 we discussed that the decay rate of a particle of astrophysical origin due to gravitational CR is expected to be of the same order as the time of propagation, which implies that the energy loss rate has to be taken into account in order to estimate the LIVPs of both the particle $a$ and the theory of gravity. Computing the integral in Eq. (38) and solving the differential Eq. would give the maximum time travel $t$ possible for a given momentum $p$ in terms of $\mathbb{A}$, $\mathbb{B}$ Moore_2001 ; Kimura_2012 . Thus, constraints on $\mathbb{A}$ and $\mathbb{B}$ can be imposed using high energy astrophysical observations if the distance traveled $ct$ and the momentum $p$ are known.
@@
To illustrate how constraints can be estimated, let us consider the simple case where $\alpha=\beta=0$. Attending to Eqs. (17) and (18) we see that the process is allowed when $\mathbb{B}\geq\mathbb{A}$. Note that this condition is equivalent to the particle $a$ having a greater phase velocity than the particle $b$, $c_{a}>c_{b}$. Integration of Eq. (38) in the massless limit is straightforward and yields
@@
Here we see that constraints are more stringent for high energy particles that have traveled a long distance $ct$; note, however, that the energy of the particle plays a more relevant role since the dependence is of the form $(tp^{3})^{-1/2}$. Another interesting feature is that the constraint is only imposed for the difference between the LIVPs, so one of them remains unfixed: as long as this difference is small enough, the values of $\mathbb{A}$ and $\mathbb{B}$ can be arbitrarily large. One should be careful with this last statement as our results have been derived assuming that $|\mathbb{A}p^{\alpha}|$ and $|\mathbb{B}k^{\beta}|$ are very small compared to unity. We have checked that both of these features, the dependence on $p$ and $t$ and the impossibility of fixing both $\mathbb{A}$ and $\mathbb{B}$ simultaneously, are present for different values of $\alpha$ and $\beta$.
It is also worth mentioning the case where $\mathbb{A}=0$ and $\beta$ is a positive real number. Performing a similar calculation in the massless limit $m=0$ shows that damping from gravitational CR is not relevant if
@@
which corresponds to the result derived in Ref. Kiyota_2015 . The constraint on $\mathbb{B}^{(\beta)}p^{\beta}$ exhibits the same dependence on the energy and the time travel as Eq. (41), and we see that the value of $\mathbb{B}^{(\beta)}$ decreases for higher values of $\beta$. This has been used to derive much more stringent constraints in the gravitational sector using the absence of vacuum CR for $\beta\gtrsim 0$ rather than using direct detection of GWs, in particular the events GW150914 and GW151226 Yunes_2016 . The same behavior is observed when the dispersion relation of $a$ particles is modified, that is, when $\mathbb{A}\neq 0$.